in a mixture we know that we want 1% of a product that is made up of 50% of a particular ingredient. If another product were substituted that had 75% of that ingredient, what would we reduce the 1% of the new product to.
Answer by Boreal(15235)

Suppose the entire product weighs 1000 pounds, so the 1% additive of a product weighs 10 pounds, with 5 pounds of ingredient X.
Now, suppose 5 pounds of ingredient X is 75% of the additive. How much does the additive weigh?
5=(3/4)x
multiply both sides by 4/3
x=20/3 pounds, or 6 2/3 pounds.
Instead of 10 pounds we now need 6 2/3 pounds. And that is 20/3/1000=20/3000 of the entire product or 2/300 or 0.6667%, reduced from 1.000%.
------------------------
.01*0.50=x*0.75
.005=0.75x
x=0.006667 or 0.6667%